OpenStudy (anonymous):

The formula for converting from Celcius to Farenheit temperatures is F=9/5C+32. a. find the inverse of the formula. Is the inverse a function? b. use the inverse to find celsius temperature that corresponds to 19 degrees farenheit.

myininaya (myininaya):

(a) just solve for C

OpenStudy (anonymous):

so is the answer \[f-32\div9/5\]

myininaya (myininaya):

F-32=9C/5 5(F-32)/9=C C=5(F-32)/9

myininaya (myininaya):

b. to find C just plug in 19 where you see the F